# Inference for Heterogeneous Treatment Effects with Efficient Instruments and Machine Learning

### v1

@@ D.3 Proof of Theorem 9
$\textstyle\frac{\sqrt{Nh}(\hat{\beta}(a)-\beta(a))}{\sigma(a)}$
$\textstyle=\frac{\frac{1}{\sqrt{Nh}}\sum_{k=1}^{K}\sum_{i\in I_{k}}(Y_{i}-\hat{l}^{k}(X_{i})-\beta(a)(D_{i}-\hat{\phi}^{k}(X_{i})))(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}^{k}(X_{i}))K\left(\frac{A_{i}-a}{h}\right)}{\frac{\sigma(a)}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}(D_{i}-\hat{\phi}^{k}(X_{i}))(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}^{k}(X_{i}))K\left(\frac{A_{i}-a}{h}\right)}$
$\textstyle=\frac{\sqrt{Nh}\hat{Q}(\beta(a),a)/\sigma_{Q}(a)}{\frac{\sigma(a)/\sigma_{Q}(a)}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}(D_{i}-\hat{\phi}^{k}(X_{i}))(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}^{k}(X_{i}))K\left(\frac{A_{i}-a}{h}\right)}$
@@
$\frac{\sigma(a)/\sigma_{Q}(a)}{Nh}\sum_{i=1}^{N}(D_{i}-\phi(X_{i}))(f(Z_{i},X_{i})-\phi(X_{i}))K\left(\frac{A_{i}-a}{h}\right)-1=o_{P}(1).$
@@
$\textstyle\frac{\sigma(a)/\sigma_{Q}(a)}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}(D_{i}-\hat{\phi}^{k}(X_{i}))(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}^{k}(X_{i}))K\left(\frac{A_{i}-a}{h}\right)$
$\textstyle-\frac{\sigma(a)/\sigma_{Q}(a)}{Nh}\sum_{i=1}^{N}(D_{i}-\phi(X_{i}))(f(Z_{i},X_{i})-\phi(X_{i}))K\left(\frac{A_{i}-a}{h}\right)$ $\textstyle=o_{P}(1).$
@@ D.3.1 Proof of Lemma 22
$\frac{\sigma(a)}{\sigma_{Q}(a)}$ $=\left(\mathbb{E}[(D_{i}-\phi(X_{i}))(f(Z_{i},X_{i})-\phi(X_{i}))|A_{i}=a]p_{A}(a)\right)^{-1}$
$=\left(\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|A_{i}=a]p_{A}(a)\right)^{-1}.$ (54)
@@
$\mathbb{E}\left[(D_{i}-\phi(X_{i}))(f(Z_{i},X_{i})-\phi(X_{i}))K\left(\frac{A_{i}-a}{h}\right)\right]$
$=\mathbb{E}\left[\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|A_{i}]K\left(\frac{A_{i}-a}{h}\right)\right]$
$=h\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|A_{i}=a]p_{A}(a)+\sigma_{Q}^{2}(a)O(h^{3}).$
@@
$V_{i}=\frac{\sigma(a)/\sigma_{Q}(a)}{h}(D_{i}-\phi(X_{i}))(f(Z_{i},X_{i})-\phi(X_{i}))K\left(\frac{A_{i}-a}{h}\right)$
satisfy $\mathbb{E}[V_{i}]=1+O(h^{2})=1+o(1)$. To prove Lemma 22, it therefore remains to prove $\frac{1}{N}\sum_{i=1}^{N}(V_{i}-\mathbb{E}[V_{i}])=o_{P}(1)$. For this, using Chebyshev’s inequality, it is enough to show
$\frac{\mathbb{E}[V_{i}^{2}]}{N}=o(1).$ (55)
We calculate $\frac{\mathbb{E}[V_{i}^{2}]}{N}$ using (54), apply Lemma 24 and assertions 10, 3, 2 and 1 of Assumption 6 to get
$\textstyle\frac{\mathbb{E}[V_{i}^{2}]}{N}$ $=\textstyle\frac{\mathbb{E}\left[(D_{i}-\phi(X_{i}))^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}\right]}{Nh^{2}\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|A_{i}=a]^{2}p_{A}(a)^{2}}$
$\textstyle=\frac{h\mathbb{E}\left[(D_{i}-\phi(X_{i}))^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|A_{i}=a\right]p_{A}(a)\int_{-\infty}^{\infty}K(x)^{2}\mathrm{d}x+\sigma_{Q}^{2}(a)O(h^{3})}{Nh^{2}\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|A_{i}=a]^{2}p_{A}(a)^{2}}$
$\textstyle\lesssim\frac{h\sigma_{Q}^{2}(a)+h^{3}\sigma_{Q}^{2}(a)}{Nh^{2}\sigma_{Q}^{4}(a)}$
$\textstyle\lesssim\frac{1}{Nh\sigma_{Q}^{2}(a)}.$
Since by the assumption of Theorem 9, $\sigma_{Q}^{2}(a)\gg(Nh)^{-1}$, (55) follows. This concludes the proof of Lemma 22.
@@ D.3.2 Proof of Lemma 23
Let $\Delta\phi_{i}^{k}=\phi(X_{i})-\hat{\phi}^{k}(X_{i})$ and $\Delta f_{i}^{k}=f(X_{i})-\hat{f}^{k}(X_{i})$ as before. Then, the quantity given in Lemma 23 is equal to
$\frac{\sigma(a)/\sigma_{Q}(a)}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}\bigl{[}(D_{i}-\phi(X_{i}))(\Delta\phi_{i}^{k}-\Delta f_{i}^{k})$
$\quad\quad\quad+\Delta\phi_{i}^{k}(f(Z_{i},X_{i})-\phi(X_{i}))$
$\quad\quad\quad+\Delta\phi_{i}^{k}(\Delta\phi_{i}^{k}-\Delta f_{i}^{i})\bigr{]}K\left(\frac{A_{i}-a}{h}\right).$
Moreover, we know from (54) and assertions 1 and 2 of Assumption 6 that $\sigma(a)/\sigma_{Q}(a)\asymp\frac{1}{\sigma_{Q}^{2}(a)}$ Hence, it remains to show that for $k=1,\ldots,K$,
$\frac{1}{nh\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}(D_{i}-\phi(X_{i}))\Delta f_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)=o_{P}(1)$ (56)
$\frac{1}{nh\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}(D_{i}-\phi(X_{i}))\Delta\phi_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)=o_{P}(1)$ (57)
$\frac{1}{nh\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))\Delta\phi_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)=o_{P}(1)$ (58)
$\frac{1}{nh\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}\Delta f_{i}^{k}\Delta\phi_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)=o_{P}(1)$ (59)
$\frac{1}{nh\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}(\Delta\phi_{i}^{k})^{2}K\left(\frac{A_{i}-a}{h}\right)=o_{P}(1).$ (60)
@@
$\left|\frac{1}{nh\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}(D_{i}-\phi(X_{i}))\Delta f_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)\right|$
$\leq\frac{1}{\sigma_{Q}^{2}(a)}{\sqrt{\frac{1}{nh}\sum_{i\in I_{k}}(D_{i}-\phi(X_{i}))^{2}K\left(\frac{A_{i}-a}{h}\right)}}{\sqrt{\frac{1}{nh}\sum_{i\in I_{k}}(\Delta f_{i}^{k})^{2}K\left(\frac{A_{i}-a}{h}\right)}}.$
@@
$\frac{1}{nh}\sum_{i\in I_{k}}(D_{i}-\phi(X_{i}))^{2}K\left(\frac{A_{i}-a}{h}\right)$ $=O_{P}(1)$ (61)
$\frac{1}{\sigma_{Q}^{4}(a)nh}\sum_{i\in I_{k}}(\Delta f_{i}^{k})^{2}K\left(\frac{A_{i}-a}{h}\right)$ $=o_{P}(1).$ (62)
(61) can be obtained as in the proof of Lemma 20 using Assertions 11 and 14 of Assumption 6 and Lemma 24. (62) follows similarly to the proof of (37) by noticing that $\frac{\|K\|_{\infty}}{\sigma_{Q}(a)^{4}h}\|f-\hat{f}^{k}\|_{L_{2}}^{2}=o_{P}(1)$ by the assumption in Theorem 9. Hence, we obtain (56). In the same way, one can obtain (57). For (58), it follows from (36) that
$\frac{1}{nh\sigma_{Q}^{2}(a)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))\Delta\phi_{i}^{k}K\left(\frac{A_{i}-a}{h}\right)=\frac{1}{\sigma_{Q}(a)\sqrt{nh}}o_{P}(1).$
By the assumption of Theorem 9, $\sigma_{Q}^{2}(a)\gg(Nh)^{-1}$ and hence, (58) follows. In the same way, (59) and (60) follow from (40) and (39).
@@ D.4 Proof of Theorem 10
We know from Lemma 22 and Lemma 23 that for the denominator of $\hat{\sigma}^{2}(a)$ defined in (24),
$\left(\frac{\sigma(a)/\sigma_{Q}(a)}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}(D_{i}-\hat{\phi}^{k}(X_{i}))(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}^{k}(X_{i}))K\left(\frac{A_{i}-a}{h}\right)\right)^{2}$
@@
$\frac{1}{Nh\sigma_{Q}^{2}(a)}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(Y_{i}-\hat{l}^{k}(X_{i})-\hat{\beta}(a)(D_{i}-\hat{\phi}^{k}(X_{i}))\right)^{2}\cdot$
$\qquad\qquad\qquad\qquad\cdot\left(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}^{k}(X_{i}))\right)^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$
@@
This equality was shown to hold if one replaces $\hat{\beta}(a)$ by $\beta(a)$ in Lemma 20 and Lemma 21. Hence, it is enough to show
$\frac{1}{Nh\sigma_{Q}^{2}(a)}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(Y_{i}-\hat{l}^{k}(X_{i})-\hat{\beta}(a)(D_{i}-\hat{\phi}^{k}(X_{i}))\right)^{2}\cdot$ (63)
$\qquad\qquad\qquad\qquad\cdot\left(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}^{k}(X_{i}))\right)^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$
$-\frac{1}{Nh\sigma_{Q}^{2}(a)}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(Y_{i}-\hat{l}^{k}(X_{i})-\beta(a)(D_{i}-\hat{\phi}^{k}(X_{i}))\right)^{2}\cdot$ (64)
$\qquad\qquad\qquad\qquad\quad\cdot\left(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}^{k}(X_{i}))\right)^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$
@@
$\Xi_{i}^{k}$ $=\frac{1}{\sqrt{h}\sigma_{Q}(a)}(\beta(a)-\hat{\beta}(a))(D_{i}-\hat{\phi}^{k}(X_{i}))\cdot$
$\qquad\qquad\qquad\cdot\left(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}^{k}(X_{i})\right)K\left(\frac{A_{i}-a}{h}\right)$
$\Theta_{i}^{k}$ $=\frac{1}{\sqrt{h}\sigma_{Q}(a)}\left(Y_{i}-\hat{l}^{k}(X_{i})-\beta(a)(D_{i}-\hat{\phi}^{k}(X_{i}))\right)\cdot$
$\qquad\qquad\qquad\cdot\left(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}^{k}(X_{i})\right)K\left(\frac{A_{i}-a}{h}\right).$
@@
$\frac{1}{N}\sum_{i=1}^{N}(\Xi_{i}^{k})^{2}$ $=\frac{(\beta(a)-\hat{\beta}(a))^{2}}{Nh\sigma_{Q}^{2}(a)}\sum_{k=1}^{K}\sum_{i\in I_{k}}(D_{i}-\hat{\phi}_{k}(X_{i}))^{2}\cdot$
$\qquad\qquad\qquad\qquad\qquad\quad\cdot\left(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}^{k}(X_{i})\right)^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}.$
By Theorem 9, we know that $\sqrt{Nh}(\hat{\beta}(a)-\beta(a))/\sigma(a)=O_{P}(1)$. From (54) and assertions 1 and 2 of Assumption 6, we know that $\sigma(a)\asymp 1/\sigma_{Q}(a)$. Since by assumption $\sigma_{Q}(a)\gg(Nh)^{-1/2}$, it follows that $(\hat{\beta}(a)-\beta(a))^{2}=o_{P}(1)$. It remains to prove that
$\frac{1}{Nh\sigma_{Q}^{2}(a)}\sum_{k=1}^{K}\sum_{i\in I_{k}}(D_{i}-\hat{\phi}^{k}(X_{i}))^{2}\left(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}^{k}(X_{i})\right)^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$
@@
As before, we first consider the “oracle” version of (68) with $\hat{\phi}^{k}$ and $\hat{f}^{k}$ replaced by $\phi$ and $f$. For this, note that by assertion 10 of Assumption 6 and Lemma 24,
$\mathbb{E}\left[\frac{1}{Nh\sigma_{Q}^{2}(a)}\sum_{i=1}^{N}(D_{i}-\phi(X_{i}))^{2}\left(f(Z_{i},X_{i})-\phi(X_{i})\right)^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}\right]$ $=O(1)$

### v2

@@ D.3 Proof of Theorem 8
$\textstyle\frac{\sqrt{Nh}(\hat{\beta}(v)-\beta(v))}{\sigma(v)}$
$\textstyle=\frac{\frac{1}{\sqrt{Nh}}\sum_{k=1}^{K}\sum_{i\in I_{k}}(Y_{i}-\hat{l}^{k}(X_{i})-\beta(v)(D_{i}-\hat{\phi}_{1}^{k}(X_{i})))(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i}))K\left(\frac{V_{i}-v}{h}\right)}{\frac{\sigma(v)}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}(D_{i}-\hat{\phi}_{1}^{k}(X_{i}))(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i}))K\left(\frac{V_{i}-v}{h}\right)}$
$\textstyle=\frac{\sqrt{Nh}\hat{Q}(\beta(v),v)/\sigma_{Q}(v)}{\frac{\sigma(v)/\sigma_{Q}(v)}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}(D_{i}-\hat{\phi}_{1}^{k}(X_{i}))(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i}))K\left(\frac{V_{i}-v}{h}\right)}$
@@
$\frac{\sigma(v)/\sigma_{Q}(v)}{Nh}\sum_{i=1}^{N}(D_{i}-\phi(X_{i}))(f(Z_{i},X_{i})-\phi(X_{i}))K\left(\frac{V_{i}-v}{h}\right)-1=o_{P}(1).$
@@
$\textstyle\frac{\sigma(v)/\sigma_{Q}(v)}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}(D_{i}-\hat{\phi}_{1}^{k}(X_{i}))(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i}))K\left(\frac{V_{i}-v}{h}\right)$
$\textstyle-\frac{\sigma(v)/\sigma_{Q}(v)}{Nh}\sum_{i=1}^{N}(D_{i}-\phi(X_{i}))(f(Z_{i},X_{i})-\phi(X_{i}))K\left(\frac{V_{i}-v}{h}\right)$ $\textstyle=o_{P}(1).$
@@ D.3.1 Proof of Lemma 21
$\frac{\sigma(v)}{\sigma_{Q}(v)}$ $=\left(\mathbb{E}[(D_{i}-\phi(X_{i}))(f(Z_{i},X_{i})-\phi(X_{i}))|V_{i}=v]p_{V}(v)\right)^{-1}$
$=\left(\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=v]p_{V}(v)\right)^{-1}.$ (54)
@@
$\mathbb{E}\left[(D_{i}-\phi(X_{i}))(f(Z_{i},X_{i})-\phi(X_{i}))K\left(\frac{V_{i}-v}{h}\right)\right]$
$=\mathbb{E}\left[\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}]K\left(\frac{V_{i}-v}{h}\right)\right]$
$=h\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=v]p_{V}(v)+\sigma_{Q}^{2}(v)O(h^{3}).$
@@
$U_{i}=\frac{\sigma(v)/\sigma_{Q}(v)}{h}(D_{i}-\phi(X_{i}))(f(Z_{i},X_{i})-\phi(X_{i}))K\left(\frac{V_{i}-v}{h}\right)$
satisfy $\mathbb{E}[U_{i}]=1+O(h^{2})=1+o(1)$. To prove Lemma 21, it therefore remains to prove $\frac{1}{N}\sum_{i=1}^{N}(U_{i}-\mathbb{E}[U_{i}])=o_{P}(1)$. For this, using Chebyshev’s inequality, it is enough to show
$\frac{\mathbb{E}[U_{i}^{2}]}{N}=o(1).$ (55)
We calculate $\frac{\mathbb{E}[U_{i}^{2}]}{N}$ using (54), apply Lemma 23 and assertions 10, 3, 2 and 1 of Assumption 6 to get
$\textstyle\frac{\mathbb{E}[U_{i}^{2}]}{N}$ $=\textstyle\frac{\mathbb{E}\left[(D_{i}-\phi(X_{i}))^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}\right]}{Nh^{2}\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=v]^{2}p_{V}(v)^{2}}$
$\textstyle=\frac{h\mathbb{E}\left[(D_{i}-\phi(X_{i}))^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=v\right]p_{V}(v)\int_{-\infty}^{\infty}K(x)^{2}\mathrm{d}x+\sigma_{Q}^{2}(v)O(h^{3})}{Nh^{2}\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=v]^{2}p_{V}(v)^{2}}$
$\textstyle\lesssim\frac{h\sigma_{Q}^{2}(v)+h^{3}\sigma_{Q}^{2}(v)}{Nh^{2}\sigma_{Q}^{4}(v)}$
$\textstyle\lesssim\frac{1}{Nh\sigma_{Q}^{2}(v)}.$
Since by the assumption of Theorem 8, $\sigma_{Q}^{2}(v)\gg(Nh)^{-1}$, (55) follows. This concludes the proof of Lemma 21.
@@ D.3.2 Proof of Lemma 22
Let $\Delta\phi_{1,i}^{k}=\phi(X_{i})-\hat{\phi}_{1}^{k}(X_{i})$, $\Delta\phi_{2,i}^{k}=\phi(X_{i})-\hat{\phi}_{2}^{k}(X_{i})$ and $\Delta f_{i}^{k}=f(X_{i})-\hat{f}^{k}(X_{i})$ as before. Then, the quantity given in Lemma 22 is equal to
$\frac{\sigma(v)/\sigma_{Q}(v)}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}\bigl[(D_{i}-\phi(X_{i}))(\Delta\phi_{2,i}^{k}-\Delta f_{i}^{k})$
$\quad\quad\quad+\Delta\phi_{1,i}^{k}(f(Z_{i},X_{i})-\phi(X_{i}))$
$\quad\quad\quad+\Delta\phi_{1,i}^{k}(\Delta\phi_{2,i}^{k}-\Delta f_{i}^{i})\bigr]K\left(\frac{V_{i}-v}{h}\right).$
Moreover, we know from (54) and assertions 1 and 2 of Assumption 6 that $\sigma(v)/\sigma_{Q}(v)\asymp\frac{1}{\sigma_{Q}^{2}(v)}$ Hence, it remains to show that for $k=1,\ldots,K$,
$\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(D_{i}-\phi(X_{i}))\Delta f_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)=o_{P}(1)$ (56)
$\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(D_{i}-\phi(X_{i}))\Delta\phi_{2,i}^{k}K\left(\frac{V_{i}-v}{h}\right)=o_{P}(1)$ (57)
$\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))\Delta\phi_{1,i}^{k}K\left(\frac{V_{i}-v}{h}\right)=o_{P}(1)$ (58)
$\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}\Delta f_{i}^{k}\Delta\phi_{1,i}^{k}K\left(\frac{V_{i}-v}{h}\right)=o_{P}(1)$ (59)
$\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}\Delta\phi_{1,i}^{k}\Delta\phi_{2,i}^{k}K\left(\frac{V_{i}-v}{h}\right)=o_{P}(1).$ (60)
@@
$\left|\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(D_{i}-\phi(X_{i}))\Delta f_{i}^{k}K\left(\frac{V_{i}-v}{h}\right)\right|$
$\leq\frac{1}{\sigma_{Q}^{2}(v)}{\sqrt{\frac{1}{nh}\sum_{i\in I_{k}}(D_{i}-\phi(X_{i}))^{2}K\left(\frac{V_{i}-v}{h}\right)}}{\sqrt{\frac{1}{nh}\sum_{i\in I_{k}}(\Delta f_{i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)}}.$
@@
$\frac{1}{nh}\sum_{i\in I_{k}}(D_{i}-\phi(X_{i}))^{2}K\left(\frac{V_{i}-v}{h}\right)$ $=O_{P}(1)$ (61)
$\frac{1}{\sigma_{Q}^{4}(v)nh}\sum_{i\in I_{k}}(\Delta f_{i}^{k})^{2}K\left(\frac{V_{i}-v}{h}\right)$ $=o_{P}(1).$ (62)
(61) can be obtained as in the proof of Lemma 19 using Assertions 11 and 14 of Assumption 6 and Lemma 23. (62) follows similarly to the proof of (37) by noticing that $\frac{\|K\|_{\infty}}{\sigma_{Q}^{4}(v)h}\|f-\hat{f}^{k}\|_{L_{2}}^{2}=o_{P}(1)$ by the assumption in Theorem 8. Hence, we obtain (56). In the same way, one can obtain (57). For (58), it follows from (36) that
$\frac{1}{nh\sigma_{Q}^{2}(v)}\sum_{i\in I_{k}}(f(Z_{i},X_{i})-\phi(X_{i}))\Delta\phi_{1,i}^{k}K\left(\frac{V_{i}-v}{h}\right)=\frac{1}{\sigma_{Q}(v)\sqrt{nh}}o_{P}(1).$
By the assumption of Theorem 8, $\sigma_{Q}^{2}(v)\gg(Nh)^{-1}$ and hence, (58) follows. In the same way, (59) and (60) follow from (40) and (39).
@@ D.4 Proof of Theorem 9
We know from Lemma 21 and Lemma 22 that for the denominator of $\hat{\sigma}^{2}(v)$ defined in (24),
$\left(\frac{\sigma(v)/\sigma_{Q}(v)}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}(D_{i}-\hat{\phi}_{1}^{k}(X_{i}))(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i}))K\left(\frac{V_{i}-v}{h}\right)\right)^{2}$
@@
$\frac{1}{Nh\sigma_{Q}^{2}(v)}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(Y_{i}-\hat{l}^{k}(X_{i})-\hat{\beta}(v)(D_{i}-\hat{\phi}_{1}^{k}(X_{i}))\right)^{2}\cdot$
$\qquad\qquad\qquad\qquad\cdot\left(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i}))\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$
@@
This equality was shown to hold if one replaces $\hat{\beta}(v)$ by $\beta(v)$ in Lemma 19 and Lemma 20. Hence, it is enough to show
$\frac{1}{Nh\sigma_{Q}^{2}(v)}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(Y_{i}-\hat{l}^{k}(X_{i})-\hat{\beta}(v)(D_{i}-\hat{\phi}_{1}^{k}(X_{i}))\right)^{2}\cdot$ (63)
$\qquad\qquad\qquad\qquad\cdot\left(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i}))\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$
$-\frac{1}{Nh\sigma_{Q}^{2}(v)}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(Y_{i}-\hat{l}^{k}(X_{i})-\beta(v)(D_{i}-\hat{\phi}_{1}^{k}(X_{i}))\right)^{2}\cdot$ (64)
$\qquad\qquad\qquad\qquad\quad\cdot\left(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i}))\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$
@@
$\Xi_{i}^{k}$ $=\frac{1}{\sqrt{h}\sigma_{Q}(v)}(\beta(v)-\hat{\beta}(v))(D_{i}-\hat{\phi}_{1}^{k}(X_{i}))\cdot$
$\qquad\qquad\qquad\cdot\left(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i})\right)K\left(\frac{V_{i}-v}{h}\right)$
$\Theta_{i}^{k}$ $=\frac{1}{\sqrt{h}\sigma_{Q}(v)}\left(Y_{i}-\hat{l}^{k}(X_{i})-\beta(v)(D_{i}-\hat{\phi}_{1}^{k}(X_{i}))\right)\cdot$
$\qquad\qquad\qquad\cdot\left(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i})\right)K\left(\frac{V_{i}-v}{h}\right).$
@@
$\frac{1}{N}\sum_{i=1}^{N}(\Xi_{i}^{k})^{2}$ $=\frac{(\beta(v)-\hat{\beta}(v))^{2}}{Nh\sigma_{Q}^{2}(v)}\sum_{k=1}^{K}\sum_{i\in I_{k}}(D_{i}-\hat{\phi}_{1}^{k}(X_{i}))^{2}\cdot$
$\qquad\qquad\qquad\qquad\qquad\quad\cdot\left(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i})\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}.$
By Theorem 8, we know that $\sqrt{Nh}(\hat{\beta}(v)-\beta(v))/\sigma(v)=O_{P}(1)$. From (54) and assertions 1 and 2 of Assumption 6, we know that $\sigma(v)\asymp 1/\sigma_{Q}(v)$. Since by assumption $\sigma_{Q}(v)\gg(Nh)^{-1/2}$, it follows that $(\hat{\beta}(v)-\beta(v))^{2}=o_{P}(1)$. It remains to prove that
$\frac{1}{Nh\sigma_{Q}^{2}(v)}\sum_{k=1}^{K}\sum_{i\in I_{k}}(D_{i}-\hat{\phi}_{1}^{k}(X_{i}))^{2}\left(\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i})\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$
@@
As before, we first consider the “oracle” version of (68) with $\hat{\phi}_{1,2}^{k}$ and $\hat{f}^{k}$ replaced by $\phi$ and $f$. For this, note that by assertion 10 of Assumption 6 and Lemma 23,
$\mathbb{E}\left[\frac{1}{Nh\sigma_{Q}^{2}(v)}\sum_{i=1}^{N}(D_{i}-\phi(X_{i}))^{2}\left(f(Z_{i},X_{i})-\phi(X_{i})\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}\right]$ $=O(1)$